# Can language models learn analogical reasoning? Investigating training objectives and comparisons to human performance

### v1

@@ Appendix B Results from other BERT architectures
OVERALL 0.51 0.97 0.06 0.52 0.85 0.19 0.66 0.83 0.48
SAT 0.50 0.98 0.02 0.51 0.76 0.26 0.53 0.76 0.29
U2 0.51 0.96 0.06 0.55 0.81 0.29 0.54 0.71 0.37
U4 0.50 0.96 0.05 0.49 0.73 0.26 0.54 0.73 0.35
SCAN 0.52 0.97 0.07 0.52 0.9 0.13 0.74 0.89 0.58
SCAN - Science 0.53 0.99 0.06 0.54 0.9 0.18 0.75 0.94 0.56
SCAN - Metaphor 0.52 0.97 0.07 0.51 0.9 0.12 0.74 0.88 0.59
@@
OVERALL 0.50 1.00 0.00 0.54 0.39 0.69 0.59 0.47 0.71
SAT 0.50 1.00 0.00 0.52 0.30 0.74 0.54 0.45 0.64
U2 0.50 1.00 0.00 0.50 0.34 0.65 0.53 0.41 0.65
U4 0.50 1.00 0.00 0.49 0.32 0.65 0.52 0.42 0.63
SCAN 0.50 1.00 0.00 0.56 0.44 0.69 0.59 0.47 0.71
SCAN - Science 0.50 1.00 0.00 0.62 0.60 0.64 0.63 0.49 0.76
SCAN - Metaphor 0.50 1.00 0.00 0.55 0.40 0.70 0.68 0.57 0.80
@@
BERT a-b non-tuned Simple Classifier BERT a-b
@@
OVERALL 0.56 0.54 0.58 0.50 0.97 0.03 0.71 0.91 0.52
SAT 0.52 0.54 0.51 0.49 0.95 0.04 0.57 0.85 0.29
U2 0.51 0.48 0.55 0.50 0.96 0.04 0.59 0.88 0.29
U4 0.53 0.55 0.52 0.50 0.94 0.05 0.58 0.88 0.29
SCAN 0.58 0.55 0.61 0.51 0.99 0.02 0.80 0.93 0.67
SCAN - Science 0.60 0.61 0.60 0.51 0.97 0.05 0.84 1.00 0.69
SCAN - Metaphor 0.58 0.54 0.62 0.50 0.99 0.02 0.79 0.92 0.66
@@
OVERALL 0.50 0.63 0.81
SAT 0.51 0.60 0.81
U2 0.48 0.62 0.83
U4 0.53 0.63 0.80
SCAN 0.49 0.63 0.81
SCAN - Science 0.49 0.66 0.84
SCAN - Metaphor 0.49 0.63 0.81
@@
OVERALL 0.53 0.66 0.68
SAT 0.55 0.65 0.64
U2 0.53 0.67 0.67
U4 0.52 0.67 0.67
SCAN 0.53 0.65 0.69
SCAN - Science 0.54 0.64 0.72
SCAN - Metaphor 0.53 0.65 0.68
@@
BERT a-b non-tuned Simple Classifier BERT a-b
OVERALL 0.54 0.60 0.88
SAT 0.53 0.59 0.86
U2 0.55 0.65 0.86
U4 0.56 0.61 0.89
SCAN 0.53 0.60 0.88
SCAN - Science 0.56 0.60 0.89
SCAN - Metaphor 0.52 0.60 0.88
@@
OVERALL 0.48 0.51 0.45 0.58 0.52 0.63 0.69 0.70 0.68
Near Overall 0.47 0.52 0.41 0.62 0.57 0.67 0.73 0.75 0.71
Categorical 0.61 0.77 0.44 0.70 0.70 0.70 0.70 0.65 0.75
Causal 0.52 0.55 0.48 0.40 0.30 0.50 0.66 0.68 0.64
Compositional 0.28 0.25 0.30 0.75 0.70 0.80 0.82 0.91 0.73
Far Overall 0.50 0.50 0.49 0.53 0.47 0.60 0.65 0.65 0.64
Categorical 0.48 0.52 0.43 0.60 0.60 0.60 0.68 0.63 0.72
Causal 0.59 0.59 0.58 0.45 0.30 0.60 0.56 0.64 0.48
Compositional 0.44 0.40 0.47 0.55 0.50 0.60 0.70 0.67 0.73
@@
Overall 0.57 0.6 0.54 0.58 0.55 0.6 0.57 0.58 0.56
Near Overall 0.59 0.59 0.59 0.63 0.57 0.70 0.61 0.60 0.61
Categorical 0.67 0.69 0.64 0.80 0.70 0.90 0.59 0.60 0.57
Causal 0.61 0.66 0.55 0.45 0.40 0.5 0.58 0.57 0.58
Compositional 0.50 0.42 0.57 0.65 0.60 0.70 0.66 0.63 0.68
Far Overall 0.55 0.60 0.49 0.52 0.53 0.50 0.53 0.55 0.51
Categorical 0.58 0.60 0.56 0.65 0.50 0.80 0.60 0.62 0.58
Causal 0.53 0.64 0.41 0.50 0.50 0.50 0.57 0.58 0.55
Compositional 0.54 0.57 0.50 0.40 0.60 0.20 0.42 0.45 0.39
@@
BERT a-b non-tuned Simple Classifier BERT a-b
@@
OVERALL 0.50 0.49 0.51 0.58 0.58 0.57 0.64 0.63 0.65
Near Overall 0.58 0.58 0.57 0.55 0.53 0.57 0.67 0.63 0.70
Categorical 0.60 0.59 0.61 0.55 0.60 0.50 0.60 0.60 0.60
Causal 0.60 0.61 0.58 0.65 0.60 0.70 0.80 0.70 0.90
Compositional 0.53 0.55 0.51 0.45 0.40 0.50 0.60 0.60 0.60
Far Overall 0.43 0.39 0.46 0.60 0.63 0.57 0.62 0.63 0.60
Categorical 0.42 0.45 0.39 0.45 0.60 0.30 0.70 0.60 0.80
Causal 0.41 0.37 0.44 0.75 0.70 0.80 0.50 0.60 0.40
Compositional 0.45 0.36 0.54 0.60 0.60 0.60 0.65 0.70 0.60

### v4

@@ Appendix D Results from other architectures: Different Models
OVERALL 0.52 0.85 0.19 0.66 0.83 0.48 0.66 0.83 0.48
SAT 0.51 0.76 0.26 0.53 0.76 0.29 0.53 0.76 0.29
U2 0.55 0.81 0.29 0.54 0.71 0.37 0.54 0.71 0.37
U4 0.49 0.73 0.26 0.54 0.73 0.35 0.54 0.73 0.35
SCAN 0.52 0.9 0.13 0.74 0.89 0.58 0.74 0.89 0.58
SCAN - Science 0.54 0.9 0.18 0.75 0.94 0.56 0.75 0.94 0.56
SCAN - Metaphor 0.51 0.9 0.12 0.74 0.88 0.59 0.74 0.88 0.59
@@
OVERALL 0.54 0.39 0.69 0.50 1.00 0.00 0.59 0.47 0.71
SAT 0.52 0.30 0.74 0.50 1.00 0.00 0.54 0.45 0.64
U2 0.50 0.34 0.65 0.50 1.00 0.00 0.53 0.41 0.65
U4 0.49 0.32 0.65 0.50 1.00 0.00 0.52 0.42 0.63
SCAN 0.56 0.44 0.69 0.50 1.00 0.00 0.63 0.49 0.76
SCAN - Science 0.62 0.60 0.64 0.50 1.00 0.00 0.68 0.57 0.80
SCAN - Metaphor 0.55 0.40 0.70 0.50 1.00 0.00 0.61 0.48 0.75
@@
RoBERTa a-b non-tuned Simple Classifier RoBERTa a-b
@@
OVERALL 0.51 0.92 0.10 0.55 0.52 0.59 0.73 0.84 0.62
SAT 0.51 0.88 0.14 0.52 0.5 0.53 0.61 0.76 0.47
U2 0.50 0.88 0.13 0.51 0.45 0.56 0.59 0.76 0.42
U4 0.50 0.87 0.12 0.54 0.52 0.57 0.59 0.77 0.42
SCAN 0.51 0.95 0.08 0.57 0.53 0.62 0.82 0.90 0.74
SCAN - Science 0.51 0.90 0.12 0.59 0.59 0.60 0.88 0.98 0.78
SCAN - Metaphor 0.51 0.96 0.07 0.57 0.52 0.62 0.80 0.88 0.73
@@
OVERALL 0.63 0.50 0.81
SAT 0.60 0.51 0.81
U2 0.62 0.48 0.83
U4 0.63 0.53 0.80
SCAN 0.63 0.49 0.81
SCAN - Science 0.66 0.49 0.84
SCAN - Metaphor 0.63 0.49 0.81
@@
OVERALL 0.66 0.53 0.68
SAT 0.65 0.55 0.64
U2 0.67 0.53 0.67
U4 0.67 0.52 0.67
SCAN 0.65 0.53 0.69
SCAN - Science 0.64 0.54 0.72
SCAN - Metaphor 0.65 0.53 0.68
@@
RoBERTa a-b non-tuned Simple Classifier RoBERTa a-b
OVERALL 0.60 0.52 0.88
SAT 0.59 0.52 0.86
U2 0.65 0.56 0.86
U4 0.61 0.53 0.89
SCAN 0.60 0.51 0.88
SCAN - Science 0.60 0.53 0.89
SCAN - Metaphor 0.60 0.50 0.88
@@
Overall 0.58 0.52 0.63 0.48 0.51 0.45 0.69 0.70 0.68
Near Overall 0.62 0.57 0.67 0.47 0.52 0.41 0.73 0.75 0.71
Categorical 0.70 0.70 0.70 0.61 0.77 0.44 0.70 0.65 0.75
Causal 0.40 0.30 0.50 0.52 0.55 0.48 0.66 0.68 0.64
Compositional 0.75 0.70 0.80 0.28 0.25 0.30 0.82 0.91 0.73
Far Overall 0.53 0.47 0.60 0.50 0.50 0.49 0.65 0.65 0.64
Categorical 0.60 0.60 0.60 0.48 0.52 0.43 0.68 0.63 0.72
Causal 0.45 0.30 0.60 0.59 0.59 0.58 0.56 0.64 0.48
Compositional 0.55 0.50 0.60 0.44 0.40 0.47 0.70 0.67 0.73
@@
Overall 0.58 0.55 0.60 0.57 0.60 0.54 0.57 0.58 0.56
Near Overall 0.63 0.57 0.70 0.59 0.59 0.59 0.61 0.60 0.61
Categorical 0.80 0.70 0.90 0.67 0.69 0.64 0.59 0.60 0.57
Causal 0.45 0.40 0.50 0.61 0.66 0.55 0.58 0.57 0.58
Compositional 0.65 0.60 0.70 0.50 0.42 0.57 0.66 0.63 0.68
Far Overall 0.52 0.53 0.50 0.55 0.60 0.49 0.53 0.55 0.51
Categorical 0.65 0.50 0.80 0.58 0.6 0.56 0.60 0.62 0.58
Causal 0.50 0.50 0.50 0.53 0.64 0.41 0.57 0.58 0.55
Compositional 0.40 0.60 0.20 0.54 0.57 0.50 0.42 0.45 0.39
@@
RoBERTa a-b non-tuned Simple Classifier RoBERTa a-b
@@
Overall 0.58 0.58 0.57 0.53 0.52 0.54 0.64 0.63 0.65
Near Overall 0.55 0.53 0.57 0.58 0.57 0.58 0.67 0.63 0.70
Categorical 0.55 0.60 0.50 0.60 0.61 0.59 0.60 0.60 0.60
Causal 0.65 0.60 0.70 0.55 0.56 0.53 0.80 0.70 0.90
Compositional 0.45 0.40 0.50 0.59 0.55 0.63 0.60 0.60 0.60
Far Overall 0.60 0.63 0.57 0.49 0.47 0.50 0.62 0.63 0.60
Categorical 0.45 0.60 0.30 0.46 0.49 0.42 0.70 0.60 0.80
Causal 0.75 0.70 0.80 0.56 0.56 0.56 0.50 0.60 0.40
Compositional 0.60 0.60 0.60 0.44 0.35 0.53 0.65 0.70 0.60